OpenStudy (anonymous):

Write the quadratic equation with a lead coefficient of 1 whose roots are -2+3i and -2-3i

OpenStudy (anonymous):

there is a real fast way to do this, but unfortunately the person who asked is off line

OpenStudy (anonymous):

you can do it in your head with ease

OpenStudy (anonymous):

@IrishBoy123 do you know the quick method?

OpenStudy (irishboy123):

no! i thought about uncompleting the square......

OpenStudy (crabbyoldgamer):

I don't know the fast way people are mentioning. I'd love to see it. But here is the regular/long way to do it. x = each of the roots. x = -2+3i and x = -2-3i. The factors of the quadratic are: (x+2-3i)(x+2+3i) Multiply them to get: \[x^2 +4x +16\] Since the lead coefficient is 1, your work is complete. This video shows a similar problem. It may help you. https://www.youtube.com/watch?v=v1xq95q_5L8